Determine for which value (s) of k will the matrix below be non-singular.


(9.1) A = 2−k −3

2 k+1



(9.2) A = 2 2 1

3 1 3

1 3 k


1
Expert's answer
2021-05-31T06:32:58-0400

A matrix is said to be singular when the determinant of that matrix is 0


(9.1)

"\begin{vmatrix}
 2-k & 3 \\ 
 2 & k+1
\end{vmatrix} =0"

"(2-k)(k+1)-6=0\\
k^2-k+4=0"

solving by this we did not get any real value if k , So we can say that for all values of k matrix will be non-singular matrix.



(9.2)

"\begin{vmatrix}
 2 & 2 & 1\\
 3 & 1
& 3 \\
1 & 3 & k \end{vmatrix} =0"

"2(k-9)-2(3k-3)+1(9-1)=0\\
-4k-4=0\\
\boxed{k=-1}"

So , we can say except k= -1, for all values of k the matrix will be non-singular matrix.